# Beyond Global Scanning: Adaptive Visual State Space Modeling for Salient Object Detection in Optical Remote Sensing Images

### v4

@@ Abstract
Salient object detection (SOD) in optical remote sensing images (ORSIs) faces numerous challenges, including significant variations in target scales and low contrast between targets and the background. Existing methods based on vision transformers (ViTs) and convolutional neural networks (CNNs) architectures aim to leverage both global and local features, but the difficulty in effectively integrating these heterogeneous features limits their overall performance. To overcome these limitations, we propose an adaptive state space context network (ASCNet), which builds upon the state space model mechanism to simultaneously capture long-range dependencies and enhance regional feature representation. Specifically, we employ the visual state space encoder to extract multi-scale features. To further achieve deep guidance and enhancement of these features, we first design a Multi-Level Context Module (MLCM). This module strengthens cross-layer interaction capabilities between features of different scales while enhancing the model’s structural perception, allowing it to distinguish between foreground and background more effectively. Then, we design the APVSS block as the decoder of ASCNet. This module integrates our proposed Dynamic Adaptive Granularity Scan (DAGS) and Granularity-aware Propagation Module (GPM). It performs adaptive patch scanning on feature maps enhanced by local perception, thereby capturing rich local region information and enhancing state space model’s local modeling capability. Extensive experimental results demonstrate that the proposed model achieves state-of-the-art performance, validating its effectiveness and superiority.
@@
[…]
[3] organization=School of Data Science, Lingnan University, city=Hong Kong, country=China
@@ 1 Introduction
Figure 2: Qualitative comparison results on challenging ORSIs scenes. From left to right are the RGB images, ground truth (GT), results produced by our method, and those generated by representative state-of-the-art (SOTA) methods. Our approach yields more accurate object localization and boundary delineation, while reducing misdetection and boundary errors.
@@
To effectively address the aforementioned challenges, we propose a novel network architecture named adaptive state space context network (ASCNet), which is specifically designed to tackle issues such as low object-background contrast and large variations in object scales in ORSI-SOD tasks. Built upon a conventional encoder–decoder architecture, ASCNet adopts a visual state space model[13, 35] as its core representation framework. The primary motivation lies in the fact that visual state space model employs a selective scanning algorithm to capture long-range dependencies while possessing good stability and scalability, which makes it possible to achieve an effective modeling balance between global context and local region details. To address the challenges of blurred object boundaries and low contrast between salient objects and backgrounds in ORSIs, we propose a Multi-Level Context Module (MLCM). This module integrates multi-scale features across different representation levels and employs a graph neural network[41] to explicitly model spatial relationships and contextual dependencies among regional features. As shown in Figure 2, by capturing cross-regional interactions and structural correlations, MLCM enables the network to better distinguish salient objects from complex backgrounds, thereby reducing segmentation errors caused by boundary ambiguity. Meanwhile, to handle the problem of multi-scale object variations and to alleviate the limitations of sequence modeling in local region representation, we introduce an Adaptive Patchwise Visual State Space (APVSS) block. This module first feeds feature maps into a Granularity-aware Propagation Module (GPM), which enhances local structural perception by introducing global semantic conditioning and promoting consistent interpretation of regional representations. Subsequently, a Dynamic Adaptive Granularity Scan (DAGS) operation is applied to further strengthen local modeling capability. Unlike conventional uniform scan strategies, DAGS partitions feature maps into resolution-aware spatial blocks and performs multi-directional scan within each block, with content-adaptive weighting to modulate the contribution of different directions for capturing local spatial dependencies. Through block-wise partitioning, direction-aware sequence construction, and integration across blocks, DAGS enhances fine-grained local feature representation while preserving global contextual awareness. This direction-adaptive modeling strategy significantly improves the network’s ability to handle scale-diverse targets and complex spatial layouts in ORSIs, ultimately leading to more accurate saliency prediction. We conduct extensive comparative experiments on two widely used benchmark datasets (i.e., ORSSD[23] and EORSSD[52]), and the experimental results demonstrate that ASCNet consistently outperforms existing state-of-the-art methods on ORSI-SOD tasks (as shown in Figure 1). Our main contributions can be summarized as follows:
@@ 3.1 Preliminaries
State Space Model: State space model (SSM)[15, 36] such as S4[14] and Mamba[13] are structured sequence architectures that are inspired by recurrent neural networks (RNNs) and CNNs and have the ability to scale linearly or nearly linearly with sequence length. These models originate from continuous systems, mapping an input sequence $x(t)\in\mathbb{R}^{L}$ to an output sequence $y(t)\in\mathbb{R}^{L}$ through a hidden state $h(t)\in\mathbb{R}^{N}$. More specifically, SSMs are described by the following continuous-time ordinary differential equations (ODEs):
@@
[…]
@@
$\overline{\mathbf{B}}$ $=(\Delta\mathbf{A})^{-1}(\exp(\Delta\mathbf{A})-\mathbf{I})\cdot\Delta\mathbf{B},$ (4)
@@ 3.4.2 Dynamic Adaptive Granularity Scan
$w_{i}^{d}=\frac{\exp\left(g_{i}^{d}\right)}{\sum_{d^{\prime}\in\{\rightarrow,\,\downarrow,\,\leftarrow,\,\uparrow\}}\exp\left(g_{i}^{d^{\prime}}\right)},\quad d\in\{\rightarrow,\,\downarrow,\,\leftarrow$$\uparrow\}$. (16)
@@
$\tilde{S}_{i}^{d}=w_{i}^{d}\cdot Scan^{d}(F_{i}),\quad d\in\{\rightarrow,\,\downarrow,\,\leftarrow$$\uparrow\}$. (17)
@@ 4.1 Experimental Setup
Figure 5: The qualitative comparison between our method and other seven SOTA approaches across various challenging scenarios.
@@ 4.2 Performance Comparisons
Figure 7: Visual comparison of qualitative results produced by models with different ablation settings.

### v5

@@ Abstract
Salient object detection (SOD) in optical remote sensing images (ORSIs) faces numerous challenges, including significant variations in target scales and low contrast between targets and the background. Existing methods based on vision transformers (ViTs) and convolutional neural networks (CNNs) architectures aim to leverage both global and local features, but the difficulty in effectively integrating these heterogeneous features limits their overall performance. To overcome these limitations, we propose an adaptive state space context network (ASCNet), which builds upon the state space model mechanism to simultaneously capture long-range dependencies and enhance regional feature representation. Specifically, we employ the visual state space encoder to extract multi-scale features. To further achieve deep guidance and enhancement of these features, we design a Multi-Level Context Module (MLCM), which module strengthens cross-layer interaction capabilities between features of different scales while enhancing the model’s structural perception, allowing it to distinguish between foreground and background more effectively. Then, we design the Adaptive Patchwise Visual State Space (APVSS) block as the decoder of ASCNet, which integrates our proposed Dynamic Adaptive Granularity Scan (DAGS) and Granularity-aware Propagation Module (GPM). It performs adaptive patch scanning on feature maps enhanced by local perception, thereby capturing rich local region information and enhancing state space model’s local modeling capability. Extensive experimental results demonstrate that the proposed model achieves state-of-the-art performance, validating its effectiveness and superiority.
@@ 1 Introduction
To effectively address the aforementioned challenges, we propose a novel network architecture named adaptive state space context network (ASCNet), which is specifically designed to tackle issues such as low object-background contrast and large variations in object scales in ORSI-SOD tasks. Inspired by recent advances in visual state space models [13, 35], our model adopts an adaptive state space modeling strategy to enhance local context aggregation and structured feature representation within an encoder–decoder architecture. The primary motivation lies in the fact that visual state space model employs a selective scanning algorithm to capture long-range dependencies while possessing good stability and scalability, which makes it possible to achieve an effective modeling balance between global context and local region details. To address the challenges of blurred object boundaries and low contrast between salient objects and backgrounds in ORSIs, we propose a Multi-Level Context Module (MLCM). This module integrates multi-scale features across different representation levels and employs a graph neural network[41] to explicitly model spatial relationships and contextual dependencies among regional features. As shown in Figure 2, by capturing cross-regional interactions and structural correlations, MLCM enables the network to better distinguish salient objects from complex backgrounds, thereby reducing segmentation errors caused by boundary ambiguity.
Figure 2: Our method delivers precise detection performance by balancing global and local modeling. From left to right are the RGB images, ground truth (GT), results produced by our method, and those generated by representative methods, namely MRBINet[22] and BCARNet[16]. Our approach yields more accurate object localization and boundary delineation, while reducing misdetection and boundary errors.
Meanwhile, to handle the problem of multi-scale object variations and to alleviate the limitations of sequence modeling in local region representation, we introduce an Adaptive Patchwise Visual State Space (APVSS) block. This module first feeds feature maps into a Granularity-aware Propagation Module (GPM), which enhances local structural perception by introducing global semantic conditioning and promoting consistent interpretation of regional representations. Subsequently, a Dynamic Adaptive Granularity Scan (DAGS) operation is applied to further strengthen local modeling capability. Unlike conventional uniform scan strategies, DAGS partitions feature maps into resolution-aware spatial blocks and performs multi-directional scan within each block, with content-adaptive weighting to modulate the contribution of different directions for capturing local spatial dependencies. Through block-wise partitioning, direction-aware sequence construction, and integration across blocks, DAGS enhances fine-grained local feature representation while preserving global contextual awareness. This direction-adaptive modeling strategy significantly improves the network’s ability to handle scale-diverse targets and complex spatial layouts in ORSIs, ultimately leading to more accurate saliency prediction. We conduct extensive comparative experiments on two widely used benchmark datasets (i.e., ORSSD[23] and EORSSD[52]), and the experimental results demonstrate that ASCNet consistently outperforms existing state-of-the-art methods on ORSI-SOD tasks (as shown in Figure 1). Our main contributions can be summarized as follows:
@@ 3.1 Preliminaries
State Space Model: State space model (SSM)[15, 36] such as S4[14] and Mamba[13] are structured sequence architectures that are inspired by recurrent neural networks (RNNs) and CNNs and have the ability to scale linearly or nearly linearly with sequence length. These models originate from continuous systems, mapping an input sequence $x(t)\in\mathbb{R}^{L}$ to an output sequence $y(t)\in\mathbb{R}^{L}$ through a hidden state $h(t)\in\mathbb{R}^{N}$. More specifically, SSMs are described by the following continuous-time ordinary differential equations (ODEs) in Eqs. (1)–(2):
@@
Discrete State Space Models: To integrate continuous-time SSMs into deep models, they need to be discretized. This involves transforming the continuous-time functions into discrete-time sequence-to-sequence mappings. For example, this is seen in models such as S4[14] and Mamba[13]. One common method is zero-order hold (ZOH). It introduces a time scale parameter $\Delta$, which converts the continuous parameters $\mathbf{A}$ and $\mathbf{B}$ into their discrete versions $\overline{\mathbf{A}}$ and $\overline{\mathbf{B}}$ in Eqs. (3)–(4).
@@
$\overline{\mathbf{B}}$ $=(\Delta\mathbf{A})^{-1}(\exp(\Delta\mathbf{A})-\mathbf{I})\cdot\Delta\mathbf{B}$ (4)
@@
Based on the discretized parameters, the discrete-time state update and output equations are given by Eqs. (5) and (6).
@@ 3.4.2 Dynamic Adaptive Granularity Scan
$w_{i}^{d}=\frac{\exp\left(g_{i}^{d}\right)}{\sum_{d^{\prime}\in\{\rightarrow,\,\downarrow,\,\leftarrow,\,\uparrow\}}\exp\left(g_{i}^{d^{\prime}}\right)},\quad d\in\{\rightarrow,\,\downarrow,\,\leftarrow$$\uparrow\}$ (16)
@@
$\tilde{S}_{i}^{d}=w_{i}^{d}\cdot Scan^{d}(F_{i}),\quad d\in\{\rightarrow,\,\downarrow,\,\leftarrow$$\uparrow\}$ (17)
@@ 4.1 Experimental Setup
Figure 5: The qualitative comparison between our method and other seven SOTA approaches across various challenging scenarios. Our method yields more accurate object localization and clearer boundaries across different scenarios. See the zoomed-in view for better visualization.
@@ 4.2 Performance Comparisons
Figure 7: Visual comparison of qualitative results produced by models with different ablation settings. The complete model achieves the best performance in terms of boundary delineation and accurate object identification.